Previous

Calendar

December

2007

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

25

JD329

26

JD330

27

JD331
mailed PAR-Lite sensor (s/n 060859) back to OSU

28

JD332

29

JD333 (cold, windy)
Site Visit (JB, JH): Details

see details for map of hydra inlet (US1, US2, UN1, UN2) locations

downloaded from the soil and tc loggers, swapped the pc card at the cr10_mrs logger

snow depths: facing north, l=9cm, middle=10cm, r=4cm

swapped 4 flasks at the west robot

checked on the span and N2 cylinders at the tower (span@350 psi, N2@850 psi)

brushed snow off of quacker's box, the soil, tc, noah and mrs logger boxes, and some other boxes on the cuff tower

confirmed the Netgear switch is in the trailer

checked the network cable connections for the tower network and the network to the world at russter2

Jia was collecting needle and phloem samples

SpB Off

30

JD334

1

JD335

2

JD336

3

JD337

4

JD338
SpB Returns

5

JD339

6

JD340

7

JD341

8

JD342

9

JD343
Site Visit (SpB): Details , Photos (17)

at MRS cr10 logger---reconnected the pcmcia external card storage (with a card).

Note, sample pump (knf UNMP30 KNI) for the profile system was showing a low flow rate.

10

JD344

11

JD345

12

JD346

13

JD347
Site Visit (JB, SpB): Details , Photos (14)

downloaded from cnr, prof, met, nctc, soil, tcs, noah, ucb, and ucan

at MET logger: started running met_06.csi (output of prop-vane data in physical units (m/s and deg, not mV)).

swapped pc cards at the mrs logger

cleaned krypton optics

changed the profiler pump; swapped the UNMP30 KNI (s/n 1/761551) for a new UNMP30 KNI

swapped N2 cylinder CC125555 @ 400 psi with CC198536 @ 1750 psi

swapped span cylinder CC240435 @ 200 psi with Scott-Marrin CC66823 @ ~450 psi

when we swapped span cylinders we noticed a leak in the cal system, possibly at the regulator or somewhere downstream of the regulator. Plan of attack for next site visit is to replace the cal system's dekabon tubing, maybe fittings and unions also, or at least check that they are tight

14

JD348

15

JD349

16

JD350
at 4:02 MST, russter2 locks up and the data system goes down.

17

JD351
Site Visit (JB): Details

at 10:48 am MST, russter2 rebooted and data system restarted.

swapped 4 flasks at the robot

rebooted russter2 @ ~1:15 pm for minor OS upgrade

18

JD352

19

JD353
SpB Off

20

JD354
at 23:16 MST, russter2 has I/O problems and the data system goes down.

21

JD355
Winter Solstice
JB goes to MA through 1/6/08
russter2 down.

22

JD356
russter2 down.

23

JD357
russter2 down.

24

JD358
CU Holiday (x-mas eve)
russter2 down.

25

JD359
CU Holiday (x-mas)
russter2 down.

26

JD360
russter2 down.

27

JD361
russter2 down.

28

JD362
SpB Returns
russter2 down.

29

JD363
Site Visit (SpB, EL-T): Details , Photos (3)

at 12:16 MST, russter2 rebooted and data system re-started.

30

JD364

31

JD365

1

JD1
CU Holiday (New Years Day)

2

JD2
Site Visit (JH, RJ): Details

at 4:02 MST, russter2 locks up and the data system goes down.

at 10:40 MST, russter2 rebooted and data system back up.

at 13:20 MST, russter2 locks up (again) and the data system goes down.

3

JD3 (warm)
Site Visit (SpB): Details , Photos (16)

at 10:15 MST, russter2 rebooted (data system not started).

at 17:00 MST, russter2 removed from trailer and brought to campus.

Prius Rental

4

JD4 (windy)
Site Visit (SpB, NT, SL): Details

at 16:30 MST, russter2 returned to trailer, rebooted, and data system re-started.

Jeep Rental

5

JD5